### v2

@@ 3.3 Properties of the L-Region
We now address the question on how to obtain $J_{i}$ that belongs to the L-region defined in (5). Prop. 2.1(a) suggests that any $J_{\mu}$ constitutes a valid option. The following result from [24] (included here for completeness) shows that $J_{i}$ can be used to encode simultaneously the Lyapunov inequality of the terminal cost and the invariance property of the terminal constraint. The cost and the constraint can be computed via optimization, such as in [25, 20], as well as via sampled data [30, 24]. To this end, we denote by $\delta_{C}:X\mapsto\{0,\infty\}$ the function that takes the value $\delta_{C}(x)=0$ if $x\in C$ and infinity otherwise.
@@
The following results show that for an element $J\in\mathcal{L}(X)$, the values of $(T^{\ell}J)(x_{k})$ are decreasing as $k$ increases, provided that the state sequence $\{x_{k}\}$ is generated under the policy $\tilde{\mu}$ defined by
@@
By the definition of $\tilde{U}(J,x)$, we have that
$(TJ)(x)=g(x,\tilde{u})+J\big{(}f(x,\tilde{u})\big{)}.$
The function $J$ belongs to the L-region so that
$g(x,\tilde{u})+(TJ)\big{(}f(x,\tilde{u})\big{)}\leq g(x,\tilde{u})+J\big{(}f(x,\tilde{u})\big{)}.$
Given that $g(x,\tilde{u})\in[0,\infty]$, we obtain the equality
@@ 4 Illustrating examples
    Parallel MPC provides better performance guarantees than that of the individual lookahead minimizations involved in the method (Example 4.2); see Prop. 3.3.
@@
$A=\begin{bmatrix}1&1\\ 0&1\end{bmatrix},\;B=\begin{bmatrix}1\\ 0.5\end{bmatrix}.$
The stage cost is given by $g(x,u)=x^{\prime}Qx+Ru^{2}+\delta_{C}(x,u)$, where prime denotes transposition, $Q$ is an identity matrix, and $R=1$. The set $C$ is $\{(x,u)\,|\,\|x\|_{\infty}\leq 5,\,|u|\leq 1\}$ ($\|x\|_{\infty}$ denotes the infinity norm of the vector $x$). We assume that four computational units are available, where $\ell_{i}=2$ for $i=1,2$ and $\ell_{i}=3$ for $i=3,4$. The functions $J_{i}(x)$ are computed in the form of $x^{\prime}K_{i}x+\delta_{S_{i}}(x)$ so that $J_{i}\in\mathcal{L}(X)$ by Prop. 3.4; see Appendix A for details.
As discussed in Prop. 3.1, there are two different approaches in implementing parallel MPC. The parallel approach given by the right-hand side of (13) makes use of multiple computing units. The alternative implementation given by the left-hand side of (13) computes directly the values of $(T^{\ell}\bar{J})(x)$. Compared with the parallel implementation, the alternative approach uses one computing unit and solves a single lookahead minimization problem involving binary variables. Our computational studies show that the parallel implementation significantly improves the computational efficiency compared with the alternative implementation. For example, Fig. 1 shows the computational time required for parallel implementation and alternative implementation for the states $\{x_{k}\}$ of trajectories starting from $x_{0}=[-5~{}2.7]^{\prime}$ and $[2.3~{}-0.6]^{\prime}$ and generated under the parallel MPC policy. We repeated the test with $100$ trials and averaged over the trials. We note that on average in this experiment the time required for parallel implementation is about $15\%$ of that needed for the alternative implementation.
Figure 1: The computational time of parallel implementation (labeled as Parallel in the figure) is compared to that of the alternative implementation with binary variables. The horizonal axis represents the index of the state along the state trajectories $\{x_{k}\}$, which are generated under the parallel MPC policy $\tilde{\mu}$. The computational time for the states are averaged over $100$ trials.
@@
(Switching System) In this example we show that the simplified variant of parallel MPC is well-suited for problems involving switching systems, where the control has both discrete and continuous elements. We consider a problem modified from [31, Section IV.A], where the dynamics are
$x_{k+1}=A_{d_{k}}x_{k}+B_{d_{k}}v_{k},$
where $u=(v,d)$ with $v\in\Re$ and $d\in\{1,2\}$, and the control constraint set is $U(x)=\Re\times\{1,2\}$ for all $x$. The dynamics are given by
@@
The stage cost is $g(x,u)=x^{\prime}Qx+Rv^{2}+\delta_{C}(x,v)$, where the parameters $Q$, $R$, and $C$ are the same as those in Example 4.1. Compared with the problem studied in [31, Section IV.A], we impose state and control constraints; in particular the term $\delta_{C}(x,v)$ is added to the stage cost. We use two computing units and we set $\ell_{i}=5$ for both units. We also introduce control constraint subsets $\bar{U}_{i}(x)=\Re\times\{i\}$, $i=1,2$. The details of $J_{i}$ are given in Appendix C. The results for simplified parallel MPC, computed for some states $x_{0}$, are summarised in Table 2. The cost function values are compared against a lower bound of $J^{*}$, which is given by $T^{8}\bar{J}_{0}$, where $\bar{J}_{0}(x)\equiv 0$ for all $x$. The difference between $J_{\tilde{\mu}}$ and $T^{8}\bar{J}_{0}$ defines an optimality gap of $J_{\tilde{\mu}}$. The values of the gap indicate that the policy $\tilde{\mu}$ is near optimal. We note that the parallel MPC scheme solves $4$ quadratic programs per time step, whereas the lower bound solves $2^{8}$ quadratic programs of a similar kind but larger size. For the computation of the lower bound, see Appendix D. Moreover, we also verified in these tests that the cost function values of parallel MPC policy are lower than that of the MPC policies defined by the individual lookahead.
@@ 5 CONCLUSIONS
We considered infinite horizon optimal control problems with deterministic dynamics. Drawing inspiration from parallel rollout, we proposed a parallel MPC method that computes approximate solutions. The proposed methods solve multiple lookahead minimization problems with different lookahead steps, and terminal costs and terminal constraints. We showed through analysis and examples that the proposed method admits a parallel implementation. In addition, we showed that parallel MPC provides better performance bounds than that given by the individual lookahead minimization involved in the method. Our computational examples illustrate that in addition to better bounds, parallel MPC provides better performance than each of the MPC schemes that correspond to the individual lookahead minimizations. Preliminary tests show that the computational savings carry to the larger scale problems.
@@
In the appendix, we provide computational details of the functions $J_{i}$ used in Examples 4.1, 4.2, and 4.3. By applying Prop. 3.4, all functions $J_{i}$ of the examples are constructed in the form of $V+\delta_{S}$. All these functions are computed offline before the realtime operation starts. We also provide a lower bound of $J^{*}$ obtained via value iteration. This bound is used in Example 4.3 to demonstrate the suboptimality of the proposed scheme. Note that alternative approaches for computing $J_{i}$ introduced in [25] can also be applied here.
@@
The sets $S_{2}$ and $S_{3}$ are computed in the same way as is for $S_{1}$, while the set $S_{4}$ is defined as
$S_{4}=\{x\in\Re^{2}\,|\,x^{\prime}K_{4}x\leq\alpha^{*}\}.$
The value $\alpha^{*}$ used for defining $S_{4}$ is the optimal value of the optimization problem
$\min_{\alpha\in\Re}$ $\quad\alpha$
$\mathrm{s.\,t.}$ $\quad x^{\prime}K_{4}x\leq\alpha,\;\forall x\text{ so that }(x,L_{4}x)\in C.$
@@
In this example, we compute the functions $J_{i}$ such that for some policy $\mu_{i}$, we have that
$g\big{(}x,\mu_{i}(x)\big{)}+J_{i}\big{(}f\big{(}x,\mu_{i}(x)\big{)}\big{)}\leq J_{i}(x)$
@@
Then we follow the procedure outlined in [23, Sections IV, V] to obtain a matrix
@@
as well as the set $S_{2}=\{x\in\Re^{2}\,|\,P_{2}x\leq e_{4}\}$, where $e^{n}\in\Re^{n}$ is an $n$-dimensional vector with all elements equal to $1$, and $P_{2}$ is given as
@@
For the third unit, the parameters are synthesized by solving a linear matrix inequality given in [23, Sections IV]. The obtained parameters are
@@
Similarly, the function $J_{3}$ takes the form $x^{\prime}K_{3}x+\delta_{S_{3}}(x)$ with
@@
Given that the open loop system $A_{1}$ and $A_{2}$ are both stable, we set $\mu_{4}(x)=0$ for all $x$. The function $J_{4}$ is given by $x^{\prime}K_{4}x+\delta_{S_{4}}(x)$, where
$K_{4}=\begin{bmatrix}1.9631&0\\ 0&1.9631\end{bmatrix},$
and the set $S_{4}=\{x\in\Re^{2}\,|\,P_{4}x\leq e_{4}\}$, where
$P_{4}^{\prime}=\begin{bmatrix}1&0&-1&0\\ 0&1&0&-1\end{bmatrix}.$
@@
In this example, we also define the functions $J_{i}$ such that for some policy $\mu_{i}$, $J_{i}(x)=J_{\mu_{i}}(x)$ for all $x$ in some set $S_{i}$. To this end, we set $L_{i}=-(B_{i}^{\prime}K_{i}B_{i}+R)^{-1}B_{i}^{\prime}K_{i}A_{i}$, $i=1,2$, where $K_{i}$ is the solution of the matrix equation
@@
The set $S_{i}$ is computed by the command invariantSet() in the MPT toolbox for the system $A_{i}+B_{i}L_{i}$ with the presence of the constraint set $C$. The function $J_{i}$ is defined as $x^{\prime}K_{i}x+\delta_{S_{i}}(x)$. This is the cost function of the policy that sets $d=i$, and $v=L_{i}x$ for $x\in S_{i}$. When applying the simplified variant at $x_{k}$ with $\ell_{i}=5$ and $J_{i}$, the computation on the $i$th unit takes the form $\big{(}T(\bar{T}^{4}_{i}J_{i})\big{)}(x_{k})$, where
$(\bar{T}_{i}J)(x)=\inf_{u\in\bar{U}_{i}(x)}\big{\{}g(x,u)+J\big{(}f(x,u)\big{)}\big{\}},$
with $\bar{U}_{i}(x)=\Re\times\{i\}$, $i=1,2$; cf. (18). This is equivalent to the following optimization problem:
@@ Appendix D A Lower Bound of Optimal Cost
Here we provide a lower bound of the optimal cost $J^{*}$, which is used in Example 4.3 to illustrate the suboptimality of our scheme. The result is well-known, see e.g. the proof to [5, Prop. 4.1.7(a)]. We include the result here for completion.
Proposition D.1.
Let the nonnegativity condition (2) hold and $\bar{J}_{0}\equiv 0$, i.e., $\bar{J}_{0}(x)=0$ for all $x$. Then the sequence $\{T^{m}\bar{J}_{0}\}$ is upper bounded by $J^{*}$, and is monotonically increasing, i.e., for $m=0,1,\dots$,
$(T^{m}\bar{J}_{0})(x)\leq J^{*}(x),\;(T^{m}\bar{J}_{0})(x)\leq(T^{m+1}\bar{J}_{0})(x),\quad\forall x\in X.$
In Example 4.3, the lower bound of $J^{*}(x_{k})$ is computed by setting $m=8$, which is the optimal value of the following optimization problem:

### v3

@@ 3.3 Properties of the L-Region
We now address the question on how to obtain $J_{i}$ that belongs to the L-region defined in (5). Prop. 2.1(a) suggests that any $J_{\mu}$ constitutes a valid option. The following result from [25] (included here for completeness) shows that $J_{i}$ can be used to encode simultaneously the Lyapunov inequality of the terminal cost and the invariance property of the terminal constraint. The cost and the constraint can be computed via optimization, as in [26, 20], as well as via sampled data [31, 25]. To this end, we denote by $\delta_{C}:X\mapsto\{0,\infty\}$ the function taking value 00 if $x\in C$ and infinity otherwise.
@@
We show next that the functions $\bar{J}_{1}$ and $\bar{J}_{2}$ in Example 3.3 belong to the L-region using Prop. 3.4.
Example 3.3.
Let us consider $\bar{J}_{1}$ and $\bar{J}_{2}$ defined in the problem (17). We define $J_{1}=V_{1}+\delta_{S_{1}}$, where $V_{1}\equiv 0$ and $S_{1}=\{0\}$. Let $\mu_{1}(0)=0$. Then one can verify that conditions (18) and (19) are satisfied with $S_{1}$, $V_{1}$, and $\mu_{1}$ in place of $S$, $V$, and $\mu$. Therefore, $J_{1}$ belongs to the L-region. Due to monotonicity of $T$, we have $\bar{J}_{1}=TJ_{1}$ belong to the L-region as well.
As for $\bar{J}_{2}$, we consider $\mu_{2}(x)=-3/2x$. In view of the control constraint $U=[-1,1]$, $\mu_{2}(x)=-3/2x$ is defined for $S_{2}=\{x\,|\,|x|\leq 2/3\}$. One can verify that $x_{2}\in S_{2}$ implies $f(x,\mu_{2}(x))\in S_{2}$. Moreover, for every $x\in S_{2}$, $J_{\mu_{2}}(x)=13/3x^{2}$. As a result, we set $V_{2}=J_{\mu_{2}}$ and $g(x,-3/2x)+V_{2}(1/2x)=V_{2}(x)$; cf. condition (19). Since $\bar{J}_{2}=V_{2}+\delta_{S_{2}}$, $\bar{J}_{2}$ belongs to the L-region.
The following result show that for an element $J\in\mathcal{L}(X)$, the values of $(T^{\ell}J)(x_{k})$ are decreasing as $k$ increases, provided that the state sequence $\{x_{k}\}$ is generated under the policy $\tilde{\mu}$ defined by
@@
By the definition of $\tilde{U}(J,x)$, we have that $(TJ)(x)=g(x,\tilde{u})+J\big{(}f(x,\tilde{u})\big{)}.$ The function $J$ belongs to the L-region so that $g(x,\tilde{u})+(TJ)\big{(}f(x,\tilde{u})\big{)}\leq g(x,\tilde{u})+J\big{(}f(x,\tilde{u})\big{)}.$ Given that $g(x,\tilde{u})\in[0,\infty]$, we obtain the equality
@@ 4 Illustrating Examples
    Parallel MPC provides better performance guarantees than that of the individual lookahead minimization involved in the method (Example 4.2); see Prop. 3.3.
@@
$A=\begin{bmatrix}1&1\\ 0&1\end{bmatrix},\qquad B=\begin{bmatrix}1\\ 0.5\end{bmatrix}.$
The stage cost is given by $g(x,u)=x^{\prime}Qx+Ru^{2}+\delta_{C}(x,u)$, where prime denotes transposition, $Q$ is an identity matrix, and $R=1$. The set $C$ is $\{(x,u)\,|\,\|x\|_{\infty}\leq 5,\,|u|\leq 1\}$ ($\|x\|_{\infty}$ denotes the infinity norm of the vector $x$). We assume that four computational units are available, where $\ell_{i}=2$ for $i=1,2$ and $\ell_{i}=3$ for $i=3,4$. The functions $J_{i}(x)$ are computed in the form of $x^{\prime}K_{i}x+\delta_{S_{i}}(x)$ so that $J_{i}\in\mathcal{L}(X)$ by Prop. 3.4; see Appendix A.
As discussed after Prop. 3.1, parallel MPC defined through Eqs. (7)-(9) can be described succinctly as the right-hand side of Eq. (14). Prop. 3.1 shows that parallel MPC is equivalent to solve directly the problem defined by left-hand side of Eq. (14), which we refer to as alternative approach from now on. Compared with the parallel MPC, this alternative approach uses one computing unit and solves a single lookahead minimization problem involving binary variables. Our computational studies show that the parallel MPC significantly improves the computational efficiency compared with the alternative approach. For example, Fig. 1 shows the computational time required for parallel MPC and alternative approach for the states $\{x_{k}\}$ of trajectories starting from $x_{0}=[-5~{}2.7]^{\prime}$ and $[2.3~{}-0.6]^{\prime}$ and generated under the parallel MPC policy. We repeated the test with $100$ trials and averaged over the trials. We note that on average in this experiment the time required for parallel MPC is about $15\%$ of that needed for the alternative approach.
Figure 1: The computational time of parallel MPC (labeled as Parallel in the figure) is compared to that of the alternative approach with binary variables. The horizontal axis represents the index of the state along the state trajectories $\{x_{k}\}$, which are generated under the parallel MPC policy $\tilde{\mu}$. The computational time is averaged over $100$ trials.
@@
(Switching System) In this example we show that the simplified variant of parallel MPC is well-suited for problems involving switching systems, where the control has both discrete and continuous elements. We consider a problem modified from [32, Section IV.A], where the dynamics are $x_{k+1}=A_{d_{k}}x_{k}+B_{d_{k}}v_{k},$ where $u=(v,d)$ with $v\in\Re$ and $d\in\{1,2\}$, and the control constraint set is $U(x)=\Re\times\{1,2\}$ for all $x$. In addition,
@@
The stage cost is $g(x,u)=x^{\prime}Qx+Rv^{2}+\delta_{C}(x,v)$, where the parameters $Q$, $R$, and $C$ are the same as those in Example 4.1. Compared with the problem studied in [32, Section IV.A], we impose state and control constraints; in particular the term $\delta_{C}(x,v)$ is added to the stage cost. We use two computing units and we set $\ell_{i}=5$ for both units. We also introduce control constraint subsets $\bar{U}_{i}(x)=\Re\times\{i\}$, $i=1,2$. The details of $J_{i}$ are given in Appendix C. The results for simplified parallel MPC, computed for some states $x_{0}$, are summarized in Table 2. The cost function values are compared against a lower bound of $J^{*}$, which is given by $T^{8}\bar{J}_{0}$, where $\bar{J}_{0}(x)\equiv 0$ for all $x$. The difference between $J_{\tilde{\mu}}$ and $T^{8}\bar{J}_{0}$ defines an optimality gap of $J_{\tilde{\mu}}$. The values of the gap indicate that the policy $\tilde{\mu}$ is near optimal. We note that the parallel MPC scheme solves $4$ quadratic programs per time step, whereas the lower bound solves $2^{8}$ quadratic programs of a similar kind but larger size. For the computation of the lower bound, see Appendix D. Moreover, we also verified in these tests that the cost function values of parallel MPC policy are lower than that of the MPC policies defined by the individual lookahead.
@@ 5 Conclusions
We considered infinite horizon optimal control problems with deterministic dynamics. Drawing inspiration from parallel rollout, we proposed a parallel MPC method that computes approximate solutions. The proposed method solves multiple lookahead minimization problems with different lookahead steps, and terminal costs and terminal constraints. We showed that the proposed method can harness the power of multiple computing units. In addition, we proved that parallel MPC provides better performance bounds than that given by the individual lookahead minimization involved in the method. Our computational examples suggest that in addition to better bounds, parallel MPC often provides better performance than each of the MPC schemes that correspond to the individual lookahead minimizations. Preliminary tests show that the computational savings carry to the larger scale problems.
@@
In the appendix, we provide computational details of the functions $J_{i}$ used in Examples 4.1, 4.2, and 4.3. Using Prop. 3.4, all functions $J_{i}$ of the examples are constructed in the form of $V+\delta_{S}$. All these functions are computed offline before the realtime operation starts. We also provide a lower bound of $J^{*}$ obtained via value iteration. This bound is used in Example 4.3 to demonstrate the suboptimality of the proposed scheme. Note that alternative approaches for computing $J_{i}$ introduced in [26] can also be applied here.
@@
The sets $S_{2}$ and $S_{3}$ are computed in the same way as is for $S_{1}$, while the set $S_{4}$ is defined as $S_{4}=\{x\in\Re^{2}\,|\,x^{\prime}K_{4}x\leq\alpha^{*}\}$. The value $\alpha^{*}$ used for defining $S_{4}$ is the optimal value of the optimization problem
$\min_{\alpha\in\Re}\ \alpha,\ \ \mathrm{s.\,t.}\ x^{\prime}K_{4}x\leq\alpha,\;\forall x\text{ so that }(x,L_{4}x)\in C.$
@@
We follow the procedure outlined in [23, Sections IV, V] and obtain
@@
as well as the set $S_{2}=\{x\in\Re^{2}\,|\,P_{2}x\leq e_{4}\}$, where $e^{n}\in\Re^{n}$ is an $n$-dimensional vector with all elements equal to $1$, and $P_{2}$ is
@@
Given that matrices $A_{1}$ and $A_{2}$ are both stable, we set $\mu_{4}(x)=0$ for all $x$. The function $J_{4}$ is given by $x^{\prime}K_{4}x+\delta_{S_{4}}(x)$, where $S_{4}=\{x\in\Re^{2}\,|\,P_{4}x\leq e_{4}\}$, and
$K_{4}=\begin{bmatrix}1.9631&0\\ 0&1.9631\end{bmatrix},\ P_{4}^{\prime}=\begin{bmatrix}1&0&-1&0\\ 0&1&0&-1\end{bmatrix}.$
@@
We define the functions $J_{i}$ such that for some policy $\mu_{i}$, $J_{i}(x)=J_{\mu_{i}}(x)$ for all $x$ in some set $S_{i}$. To this end, we set $L_{i}=-(B_{i}^{\prime}K_{i}B_{i}+R)^{-1}B_{i}^{\prime}K_{i}A_{i}$, $i=1,2$, where $K_{i}$ is the solution of the matrix equation
@@
The set $S_{i}$ is computed by the command invariantSet() in the MPT toolbox for the system $A_{i}+B_{i}L_{i}$ with the presence of the constraint set $C$. The function $J_{i}$ is defined as $x^{\prime}K_{i}x+\delta_{S_{i}}(x)$. This is the cost function of the policy that sets $d=i$, and $v=L_{i}x$ for $x\in S_{i}$. When applying the simplified variant at $x_{k}$ with $\ell_{i}=5$ and $J_{i}$, the computation on the $i$th unit takes the form $\big{(}T(\bar{T}^{4}_{i}J_{i})\big{)}(x_{k})$, where $(\bar{T}_{i}J)(x)=\inf_{u\in\bar{U}_{i}(x)}\big{\{}g(x,u)+J\big{(}f(x,u)\big{)}\big{\}},$ with $\bar{U}_{i}(x)=\Re\times\{i\}$, $i=1,2$; cf. (21). This is equivalent to the following optimization problem:
@@ Appendix D A Lower Bound of Optimal Cost
In Example 4.3, the lower bound of $J^{*}(x_{k})$ is $(T^{8}\bar{J}_{0})(x_{k})$, which is the optimal value of the following optimization problem: